OpenStudy (fernandez555):

(01.04) What is the slope of the line that passes through (3, −7) and (−1, 1)? undefined 0 −2 2

OpenStudy (the_fizicx99):

M = y2-y1/x2-x1

OpenStudy (anonymous):

your simplified answer would be 3/2.

OpenStudy (the_fizicx99):

No, 1-7/-1-3 -8/-4 = 2

OpenStudy (the_fizicx99):

@Fernandez555

OpenStudy (anonymous):

it's -2 1--7=8 -1-3=-4 8/(-4)=-2